How many moles of copper(II) sulfate will react with 16.5 grams of

aluminum?


1
Expert's answer
2021-06-11T06:44:17-0400

The balance chemical equation for the above is:


"2Al +3CuSO_4 \longrightarrow Al_2(SO_4)_3 +3Cu"


Moles of Aluminum "= \dfrac{16.5}{27} =0.6 moles"


Now by the mole mole analysis :


The moles of ,the moles of copper (ii) sulfate "= \dfrac{3}{2} \times 0.6=0.9 moles"